OpenStudy (anonymous):

my final exam (calculus 1) 5 years ago \[\lim_{n \rightarrow \infty} \sum_{k=1}^{\infty} (\sqrt{1+\frac{k}{n^2}}-1)\]

OpenStudy (anonymous):

\(\boxed{\boxed{\boxed{\boxed{\boxed{\boxed{\boxed{\boxed{\boxed{\boxed{\pi}}}}}}}}}}\)

OpenStudy (kainui):

What's your question?

OpenStudy (anonymous):

Congratulations For Golden benim yoldashim.. @mukushla

OpenStudy (anonymous):

tnx mere yaar

OpenStudy (klimenkov):

Nice problem. But I cant find the solution. I will think about and try to solve it.

OpenStudy (klimenkov):

I don't like that \(k\) has 1 degree in radicand and \(n\) has 2.

OpenStudy (anonymous):

yes nice one ...worth to try...

OpenStudy (klimenkov):

I tried to use \[\lim_{n\rightarrow \infty}\left( \frac{b-a}{n}\cdot\sum_{k=1}^{n} \ f\left(a+k\cdot\frac{b-a}{n} \right)\right)=\int_a^bf(x)\ dx\]If \(f\) is defined on \([a,b]\). \(a=0,\ b=1\). Rewrite the limit in other form: \[\lim_{n\rightarrow \infty}\left( \frac{1}{n}\cdot\sum_{k=1}^{n} \ n\cdot \left(\sqrt{1+\frac1 n \cdot\frac{k}{n}}-1\right)\right)\]Calculate the integral: \[\int_0^1\ n\cdot \left(\sqrt{1+\frac x n }-1\right)\ dx=-n + \frac23 \left(-1 + \left(1 + \frac 1n\right)^{\frac32}\right) n^2\]If we calculate the limit we will get: \[\lim_{n \rightarrow \infty} \sum_{k=1}^{\infty} \left(\sqrt{1+\frac{k}{n^2}}-1\right)=0\]

OpenStudy (klimenkov):

But I cant understand why zero.

OpenStudy (klimenkov):

I tried to calculate it numerically with Mathematica. I put k and n = 1000000 fnd it seems to be 0, but i cant understand how it can be.

OpenStudy (anonymous):

ahh... i thought it has a value lets think again...

OpenStudy (experimentx):

lol .. that would make double sum!! for n = k = 10000 mathematica seems to take forever. anyone doesn't own a supercomputer?? lol

OpenStudy (klimenkov):

Did anybody understand my solution?

OpenStudy (anonymous):

yes i workin on it...

OpenStudy (experimentx):

lol ... we aren't summing n !!

OpenStudy (anonymous):

sorry guys i lost my connection @klimenkov no problem with ur approach u just calculated lim wrong \[\lim_{n \rightarrow \infty} \left( \frac{2}{3}n^2 \left( (1+\frac{1}{n})^{\frac{3}{2}}-1 \right)-n\right)=\frac{1}{4}\]

OpenStudy (klimenkov):

We have to divide it by n and finally get zero.

OpenStudy (anonymous):

here is my solution Using Taylor series \[ (1+\frac{k}{n^2})^{\frac{1}{2}} \approx1+\frac{k}{2n^2} \] then \[\lim_{n \rightarrow \infty} \sum_{k=1}^{\infty} (\sqrt{1+\frac{k}{n^2}}-1)=\lim_{n \rightarrow \infty} \sum_{k=1}^{\infty} (1+\frac{k}{2n^2}-1)=\lim_{n \rightarrow \infty} \sum_{k=1}^{\infty} \frac{k}{2n^2}\\=\lim_{n \rightarrow \infty} \frac{n(n+1)}{4n^2}=\frac{1}{4}\]